488. The scores of 40,000 county employees who took a recent civil service paper were normally...

488.

The scores of 40,000 county employees who took a recent civil service paper were normally distributed with
μ = 70 and σ = 10.

Determine the civil service paper score (to the nearest whole number) that represents the 4th decile.
Answer #1

Z score for 4th decile = -Z0.4 = -0.25

T 11 o

x=\mu+z*\sigma

x=70-0.25*10=67.5 = 67 (ans)